OpenStudy (kpop4life123):

Calysta sells used video games. In addition to a fixed salary, she earns a commission for each video game she sells. The table shows Calysta's total earnings, y, (in dollars), from selling x video games: Video Games and Earnings Games Sold(x) Total Earnings(y) 1 78 2 86 3 94

OpenStudy (kpop4life123):

Which equation best shows the relationship between x and y? (4 points) y = x + 8 y = 8x + 70 y = 8x + 78 y = x + 78

OpenStudy (kpop4life123):

@Sean.Kirby @ash_heather1

OpenStudy (anonymous):

i know what to do i just dont know how to explain it XD

OpenStudy (kpop4life123):

lol

OpenStudy (anonymous):

just switch out the numbers where it says x in the equation put whatever number is in the x column and if it works for 1 try it on the rest of the numbers in the x colum

OpenStudy (anonymous):

there is no point in trying a because you know it can't be just +8

OpenStudy (kpop4life123):

its really hard ive been trying to solve it for two days

OpenStudy (anonymous):

okay let's try the first one

OpenStudy (anonymous):

8x+70 replace the x with 1 from the x colum so 8*1+70, what does that equal?

OpenStudy (kpop4life123):

sorry i was downstairs its 78

OpenStudy (anonymous):

there you go so since 78 is what's in the y colum that works lets try the next number

OpenStudy (anonymous):

so now you do 8*2+70

OpenStudy (kpop4life123):

86

OpenStudy (anonymous):

okay so that works

OpenStudy (anonymous):

lets try the next number

OpenStudy (anonymous):

8*3+70

OpenStudy (kpop4life123):

94

OpenStudy (anonymous):

and that one works so your anwser is B

OpenStudy (kpop4life123):

thanks!! you explained it well

OpenStudy (anonymous):

no problem :)

OpenStudy (kpop4life123):

:)